A polytropic process of air from 1034 kPa, 149 deg. C and 0.028 m3 occurs to 138 kPa in accordance with pV1.3 = C. Determine:

(a) the final temperature in deg. C

(b) the change in internal energy in kJ

(c) the work nonflow in kJ

(d) the heat in kJ from the formula using polytropic specific heat


1
Expert's answer
2021-12-02T10:08:08-0500

a)

"T_2=T_1(\frac{p_2}{p_1})^{1-\frac 1{\gamma}}=265~K=-8°C,"

b)

"A=\frac{p_1V_1}{\gamma-1}(1-(\frac{p_2}{p_1})^{1-\frac 1{\gamma}})=35.8~kPa."